you have 8 cards, 5 are green, 3 are yellow. Suppose that you randomly draw 2 cards, one at a time, WITHOUT REPLECEMENT. Let G1= 1st card is green, let G2= 2nd card is green.
find P(G1 and G2)
find P(at least one green).
find P(G2|G1)
are G2 and G1 independent events? why or why not?

Answer by Boreal(15235)

The first is (5/8)*4/7)=20/56 or 5/14.
at last one green is 1- both yellow or 1-(3/8)(2/7)=25/28
P(G2|G1) is 4/7, one less from each the numerator and the denominator.
They aren't independent because when the first one is drawn it changes the probability for the second. Sampling with replacement would make them independent.